A comet is observed to be 130 km long as it moves past an observer at a speed of 0.700c. How long does the comet seem when it travels at a speed of 0.900c with respect to the observer?
1
Expert's answer
2020-09-23T09:00:03-0400

When comet moves with the speed "v_1=0.7c", it's length:

"l_1=l_0\sqrt{1-(\frac{v_1}{c})^2}"

When it moves with the speed "v_2=0.9c":

"l_2=l_0\sqrt{1-(\frac{v_2}{c})^2}".

We have the system of equation, where we must find "l_2". Rewrite it in another form.

"l_1=l_0\sqrt{\frac{c^2-{v_1}^2}{c^2}}"

"l_2=l_0\sqrt{\frac{c^2-{v_2}^2}{c^2}}"

Divide these equations.

"\frac{l_2}{l_1}=\sqrt{\frac{c^2-{v_2}^2}{c^2-{v_1}^2}}"

Substitute numbers.

"\frac{l_2}{l_1}=\sqrt{\frac{c^2-0.81c^2}{c^2-0.49c^2}}=\sqrt{\frac{0.19}{0.51}}=0.61"

"l_2=0.61l_1=0.61*130=79" km.